Develop a current stock value for a firm that is expected to have low growth of 3% for 3 years, after which it will experience a constant-growth rate of 7%. Its required return is 14% and next year's dividend is expected to be $2.00.
Answer #1

Di=2 D2 = 2X1.03 - 2006 D3= 2.06% 1003 = 2.12 Dy = 2012 x1-03= 2.19 D5 = 2:19 x 1.07 - 2034 Terminal value (P4) = D5 g Re=14%